# Calibrating Deep Neural Network using Euclidean Distance

### v1

@@ A.8. Proof that the Focal Calibration Loss is Strictly Proper
We aim to show that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is minimized if and only if $\hat{\bm{p}}=\bm{y}$ when $\lambda>0$ and $\gamma>0$, and that it is not strictly proper when $\lambda=0$.
@@
Let $\bm{y}\in\Delta^{K}$ be the true probability vector, and $\hat{\bm{p}}\in\Delta^{K}$ be the predicted probability vector. We consider two scenarios:
(a) When $\hat{\bm{p}}=\bm{y}$.
At $\hat{\bm{p}}=\bm{y}$, the calibration loss is:
(75) $\mathcal{L}_{\text{calib}}(\bm{y},\bm{y})=\left\|\bm{y}-\bm{y}\right\|_{2}^{2}=0.$
Assuming $\bm{y}$ is a one-hot vector with $y_{j}=1$ for the true class $j$, the focal loss becomes:
(76) $\mathcal{L}_{\text{focal}}^{\gamma}(\bm{y},\bm{y})=-(1-y_{j})^{\gamma}\log y_{j}=-0^{\gamma}\cdot 0=0.$
@@
(77) $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{y},\bm{y})=0+\lambda\cdot 0=0.$
(b) When $\hat{\bm{p}}\neq\bm{y}$.
@@
(78) $\mathcal{L}_{\text{calib}}(\hat{\bm{p}},\bm{y})=\left\|\hat{\bm{p}}-\bm{y}\right\|_{2}^{2}>0.$
@@
(79) $\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\bm{p}},\bm{y})=-(1-\hat{p}_{j})^{\gamma}\log\hat{p}_{j}>0\quad\text{for }\hat{p}_{j}<1.$
@@
(80) $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\hat{\bm{p}},\bm{y})=\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\bm{p}},\bm{y})+\lambda\,\mathcal{L}_{\text{calib}}(\hat{\bm{p}},\bm{y})>0.$
@@
Since the loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ attains its unique global minimum of zero when $\hat{\bm{p}}=\bm{y}$, it is strictly proper when $\lambda>0$.
@@
(81) $\mathcal{L}_{\text{FCL}}^{\gamma,0}(\hat{\bm{p}},\bm{y})=\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\bm{p}},\bm{y}).$
@@
Consider any $\hat{\bm{p}}$ where $\hat{p}_{j}=1$ and $\hat{p}_{k}=0$ for $k\neq j$. In this case:
(82) $\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\bm{p}},\bm{y})=0.$
However, suppose $\hat{p}_{j}=q<1$ but still significantly higher than other $\hat{p}_{k}$. Depending on $\gamma$, the focal loss may not penalize this deviation strongly, and $\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\bm{p}},\bm{y})$ may still be close to zero. Therefore, the minimum loss may not be uniquely attained at $\hat{\bm{p}}=\bm{y}$.
@@
The focal loss is not strictly convex with respect to $\hat{\bm{p}}$, especially when $\gamma=0$, where it reduces to the cross-entropy loss, which is proper but not strictly proper without additional regularization.
@@
Since the minimum of $\mathcal{L}_{\text{FCL}}^{\gamma,0}$ is not uniquely achieved at $\hat{\bm{p}}=\bm{y}$, the loss is not strictly proper when $\lambda=0$.
@@
A.9. Focal Loss Relative with Entropy Regularised KL Divergence
@@
(83) $\mathcal{L}_{\text{focal}}(\hat{\bm{p}},\bm{y})$ $=-\sum_{k=1}^{K}y_{ik}(1-\hat{p}_{ik})^{\gamma}\log\hat{p}_{ik}$
@@
[…]
@@
(84) $\mathcal{L}_{\text{focal}}(\hat{\bm{p}},\bm{y})\geq KL(q\|\hat{\bm{p}})+H[q]-\gamma H[\hat{\bm{p}}]$
In the case of one-hot encoding (Delta distribution for $q$), focal loss will maximize $-\hat{p}_{k}\log\hat{p}_{k}$ (let $k$ be the ground-truth class index), the component of the entropy of $\hat{\bm{p}}$ corresponding to the ground-truth index. Thus, it will prefer learning $\hat{p}$ such that $\hat{p}_{k}$ is assigned a higher value (because of the KL term), but not too high (because of the entropy term), and will ultimately avoid preferring overconfident models (by contrast to cross-entropy loss).
Figure 7. Optimal Predicted Probability ($\hat{p_{ik}}$) for Various Values of Ground True Probability ($y_{ik}$) with Calibration Analysis. Circular markers with ”+” and ”-” signs represent the areas where the calibration term increases or decreases in visualizing how it adjust the predicted probabilities.
[…]
This plot displays the relationship between the optimal predicted probability ($\hat{p}$) and the true label probability ($q$) for a perfectly calibrated model and a miscalibrated model.
A.10. Hyperparameter Selection for Focal Calibration Loss
Table 5. $\gamma$ and $\lambda$ Selection
@@ A.10. Hyperparameter Selection for Focal Calibration Loss
A.11. Decision Boundary on Focal Loss and Focal Calibration Loss (FCL)
@@
Figure 8. Decision boundary comparison between Focal Loss (left) and Focal Calibration Loss (FCL) (right) on a synthetic dataset with hard samples. In both models, blue and red regions represent the areas classified as Class 0 and Class 1, respectively. The yellow boxes highlight regions where hard samples from both classes overlap, and the green boxes indicate areas where FCL provides a smoother decision boundary compared to Focal Loss.
[…]
The Focal Loss model attempts to correctly classify the difficult (hard) samples by creating sharp boundaries around overlapping regions. In contrast, the FCL model produces a smoother decision boundary that is not only capable of focusing on hard samples but also better calibrated, as it balances both classification accuracy and calibration.
A.12. Table 6 AdaECE Evaluation Result
Table 6. $\downarrow$ AdaECE (%) for various methods both before (pre) and after (post) applying temperature scaling.
Dataset Model Weight Decay(Guo et al., 2017) Brier Loss(Brier, 1950) MMCE(Kumar et al., 2018) Label Smooth(Szegedy et al., 2016) Focal Loss - 53(Mukhoti et al., 2020a) Dual Focal(Tao et al., 2023) Focal Calibration
@@ A.12. Table 6 AdaECE Evaluation Result
A.13. Table 7 Classwise-ECE Evaluation Result
Table 7. $\downarrow$ Classwise-ECE (%) for various methods both before (pre) and after (post) applying temperature scaling.
Dataset Model Weight Decay(Guo et al., 2017) Brier Loss(Brier, 1950) MMCE(Kumar et al., 2018) Label Smooth(Szegedy et al., 2016) Focal Loss - 53(Mukhoti et al., 2020a) Dual Focal(Tao et al., 2023) Focal Calibration
@@ A.13. Table 7 Classwise-ECE Evaluation Result
A.14. Proof of Thm. (8): Post-Processing Gap of focal calibration loss and focal loss
@@
(85) $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\kappa(\hat{\bm{p}}),\bm{y})=\mathcal{L}_{\text{focal}}(\kappa(\hat{\bm{p}}),\bm{y})+\lambda\mathcal{L}_{\text{calib}}(\kappa(\hat{\bm{p}}),\bm{y})$
@@
Following Definition 5, the post-processing gap for the FCL $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda})$is defined as:
(86) $\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\kappa(f(x)),y)]$
@@
(87) $\begin{split}\text{pGap}_{\mathcal{D}}\left(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}\right)=\ &\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{focal}}\left(f(x),y\right)\right]\\ &+\lambda\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{calib}}\left(f(x),y\right)\right]\\ &-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{focal}}\left(\kappa\left(f(x)\right),y\right)\right]\\ &-\lambda\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{calib}}\left(\kappa\left(f(x)\right),y\right)\right]\end{split}$
@@
(88) $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda=0})=\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}})$
@@
(89) $\lambda\left(\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(\kappa(f(x)),y)]\right)$
This term measures the improvement in calibration achieved through post-processing $\kappa$. The calibration loss, $\mathcal{L}_{\text{calib}}(f(x),y)$, can be defined as a squared loss: $\mathcal{L}_{\text{calib}}(\hat{\bm{p}},\bm{y})=\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}.$
@@
(90) $\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}\right]-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\kappa(\hat{p}_{ik}))^{2}\right]$
@@
[…]
@@
(92) $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda})$ $=\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}})+\lambda\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{calib}})$
@@
When $\hat{p}_{ik}>y_{ik}$, $\eta(\hat{p}_{ik})>0$. In this case, the term $2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})$ is negative, and $-\eta(\hat{p}_{ik})^{2}<0$, which means that the calibration term overall is negative. This reduction in the post-processing gap brings the prediction closer to the true label, reducing the pGap.
@@
When $\hat{p}_{ik}<y_{ik}$, $\eta(\hat{p}_{ik})>0$, and thus $2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})$ is positive. This positive term increases the predicted value, reducing the gap between $\hat{p}_{ik}$ and $y_{ik}$. Additionally, the term $-\eta(\hat{p}_{ik})^{2}$ is still negative but smaller in magnitude, ensuring that the gap is reduced.
@@
(93) $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}})\leq\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}}).$
A.15. Reliability Diagram and ROC Diagram
Figure 9. Reliability Diagram of Various Methods before Temperature Scaling(CIFAR10, ResNet-110, 300 Epochs).
[…]
tbc
Figure 10. Reliability Diagram of Various Methods before Temperature Scaling (CIFAR10, ResNet-50, 100 - 300 Epochs). The top/middle/bottom row corresponds to the results at 100/200/300 epochs. Each row shows the performance of Cross Entropy, FLSD-53, Dual Focal Loss, and Focal Calibration Loss.
[…]
tbc
(a) ResNet-50 (pre-T)
[…]
ROC curve for ResNet-50 architecture before temperature scaling, tested on SVHN dataset.
(b) ResNet-110 (pre-T)
[…]
ROC curve for ResNet-110 architecture before temperature scaling, tested on SVHN dataset.
(c) Wide-ResNet (pre-T)
[…]
ROC curve for Wide-ResNet architecture before temperature scaling, tested on SVHN dataset.
(d) DenseNet (pre-T)
[…]
ROC curve for DenseNet architecture before temperature scaling, tested on SVHN dataset.
(e) ResNet-50 (post-T)
[…]
ROC curve for ResNet-50 architecture after temperature scaling, tested on SVHN dataset.
(f) ResNet-110 (post-T)
[…]
ROC curve for ResNet-110 architecture after temperature scaling, tested on SVHN dataset.
(g) Wide-ResNet (post-T)
[…]
ROC curve for Wide-ResNet architecture after temperature scaling, tested on SVHN dataset.
(h) DenseNet (post-T)
[…]
ROC curve for DenseNet architecture after temperature scaling, tested on SVHN dataset.
Figure 11. ROC plots obtained from Different architectures trained on CIFAR-10 and tested on SVHN

### v2

@@ A.8 Proof that the Focal Calibration Loss is Strictly Proper
We aim to show that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ is minimized if and only if $\hat{\boldsymbol{p}}=\boldsymbol{y}$ when $\lambda>0$ and $\gamma>0$, and that it is not strictly proper when $\lambda=0$.
@@
Let $\boldsymbol{y}\in\Delta^{K}$ be the true probability vector, and $\hat{\boldsymbol{p}}\in\Delta^{K}$ be the predicted probability vector. We consider two scenarios:
(a) When $\hat{\boldsymbol{p}}=\boldsymbol{y}$.
At $\hat{\boldsymbol{p}}=\boldsymbol{y}$, the calibration loss is:
$\mathcal{L}_{\text{calib}}(\boldsymbol{y},\boldsymbol{y})=\left\|\boldsymbol{y}-\boldsymbol{y}\right\|_{2}^{2}=0.$ (74)
Assuming $\boldsymbol{y}$ is a one-hot vector with $y_{j}=1$ for the true class $j$, the focal loss becomes:
$\mathcal{L}_{\text{focal}}^{\gamma}(\boldsymbol{y},\boldsymbol{y})=-(1-y_{j})^{\gamma}\log y_{j}=-0^{\gamma}\cdot 0=0.$ (75)
@@
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{y},\boldsymbol{y})=0+\lambda\cdot 0=0.$ (76)
(b) When $\hat{\boldsymbol{p}}\neq\boldsymbol{y}$.
@@
$\mathcal{L}_{\text{calib}}(\hat{\boldsymbol{p}},\boldsymbol{y})=\left\|\hat{\boldsymbol{p}}-\boldsymbol{y}\right\|_{2}^{2}>0.$ (77)
@@
$\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})=-(1-\hat{p}_{j})^{\gamma}\log\hat{p}_{j}>0\quad\text{for }\hat{p}_{j}<1.$ (78)
@@
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\hat{\boldsymbol{p}},\boldsymbol{y})=\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})+\lambda\,\mathcal{L}_{\text{calib}}(\hat{\boldsymbol{p}},\boldsymbol{y})>0.$ (79)
@@
Since the loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ attains its unique global minimum of zero when $\hat{\boldsymbol{p}}=\boldsymbol{y}$, it is strictly proper when $\lambda>0$.
@@
$\mathcal{L}_{\text{FCL}}^{\gamma,0}(\hat{\boldsymbol{p}},\boldsymbol{y})=\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y}).$ (80)
@@
Consider any $\hat{\boldsymbol{p}}$ where $\hat{p}_{j}=1$ and $\hat{p}_{k}=0$ for $k\neq j$. In this case:
$\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})=0.$ (81)
However, suppose $\hat{p}_{j}=q<1$ but still significantly higher than other $\hat{p}_{k}$. Depending on $\gamma$, the focal loss may not penalize this deviation strongly, and $\mathcal{L}_{\text{focal}}^{\gamma}(\hat{\boldsymbol{p}},\boldsymbol{y})$ may still be close to zero. Therefore, the minimum loss may not be uniquely attained at $\hat{\boldsymbol{p}}=\boldsymbol{y}$.
@@
The focal loss is not strictly convex with respect to $\hat{\boldsymbol{p}}$, especially when $\gamma=0$, where it reduces to the cross-entropy loss, which is proper but not strictly proper without additional regularization.
@@
Since the minimum of $\mathcal{L}_{\text{FCL}}^{\gamma,0}$ is not uniquely achieved at $\hat{\boldsymbol{p}}=\boldsymbol{y}$, the loss is not strictly proper when $\lambda=0$.
@@
A.9 Focal Loss Relative with Entropy Regularised KL Divergence
@@
$\mathcal{L}_{\text{focal}}(\hat{\boldsymbol{p}},\boldsymbol{y})$ $=-\sum_{k=1}^{K}y_{ik}(1-\hat{p}_{ik})^{\gamma}\log\hat{p}_{ik}$ (82)
@@
$\text{By Bernoulli's inequality}\,\forall\gamma\geq 1,\text{since}\,\hat{p}_{ik}\in[0,1]$
@@
$\mathcal{L}_{\text{focal}}(\hat{\boldsymbol{p}},\boldsymbol{y})\geq KL(q\|\hat{\boldsymbol{p}})+H[q]-\gamma H[\hat{\boldsymbol{p}}$] (83)
In the case of one-hot encoding (Delta distribution for $q$), focal loss will maximize $\hat{p}_{k}\log\hat{p}_{k}$ (let $k$ be the ground-truth class index), the component of the entropy of $\hat{\boldsymbol{p}}$ corresponding to the ground-truth index. Thus, it will prefer learning $\hat{p}$ such that $\hat{p}_{k}$ is assigned a higher value (because of the KL term), but not too high (because of the entropy term), and will ultimately avoid preferring overconfident models (by contrast to cross-entropy loss).
Figure 6: Optimal Predicted Probability ($\hat{p_{ik}}$) for Various Values of Ground True Probability ($y_{ik}$) with Calibration Analysis. Circular markers with ”+” and ”-” signs represent the areas where the calibration term increases or decreases in visualizing how it adjust the predicted probabilities.
A.10 Hyperparameter Selection for Focal Calibration Loss
Table 5: $\gamma$ and $\lambda$ Selection
@@ A.10 Hyperparameter Selection for Focal Calibration Loss
A.11 Decision Boundary on Focal Loss and Focal Calibration Loss (FCL)
@@
Figure 7: Decision boundary comparison between Focal Loss (left) and Focal Calibration Loss (FCL) (right) on a synthetic dataset with hard samples. In both models, blue and red regions represent the areas classified as Class 0 and Class 1, respectively. The yellow boxes highlight regions where hard samples from both classes overlap, and the green boxes indicate areas where FCL provides a smoother decision boundary compared to Focal Loss.
A.12 Table 6 AdaECE Evaluation Result
Table 6: $\downarrow$ AdaECE (%) for various methods both before (pre) and after (post) applying temperature scaling.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
@@ A.12 Table 6 AdaECE Evaluation Result
A.13 Table 7 Classwise-ECE Evaluation Result
Table 7: $\downarrow$ Classwise-ECE (%) for various methods both before (pre) and after (post) applying temperature scaling.
Dataset Model Weight Decay (Guo et al., 2017) Brier Loss (Brier, 1950) MMCE (Kumar et al., 2018) Label Smooth (Szegedy et al., 2016) Focal Loss - 53 (Mukhoti et al., 2020a) Dual Focal (Tao et al., 2023) Focal Calibration (Ours)
@@ A.13 Table 7 Classwise-ECE Evaluation Result
A.14 Proof of Thm. (4.5): Post-Processing Gap of focal calibration loss and focal loss
@@
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\kappa(\hat{\boldsymbol{p}}),\boldsymbol{y})=\mathcal{L}_{\text{focal}}(\kappa(\hat{\boldsymbol{p}}),\boldsymbol{y})+\lambda\mathcal{L}_{\text{calib}}(\kappa(\hat{\boldsymbol{p}}),\boldsymbol{y})$ (84)
@@
Following Definition 2.3, the post-processing gap for the FCL $\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda})$is defined as:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\kappa(f(x)),y)]$ (85)
@@
$\begin{split}\text{pGap}_{\mathcal{D}}\left(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}\right)=\ &\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{focal}}\left(f(x),y\right)\right]\\ &+\lambda\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{calib}}\left(f(x),y\right)\right]\\ &-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{focal}}\left(\kappa\left(f(x)\right),y\right)\right]\\ &-\lambda\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{calib}}\left(\kappa\left(f(x)\right),y\right)\right]\end{split}$ (86)
@@
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda=0})=\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}})$ (87)
@@
$\lambda\left(\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(f(x),y)]-\mathbb{E}_{(x,y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(\kappa(f(x)),y)]\right)$ (88)
This term measures the improvement in calibration achieved through post-processing $\kappa$. The calibration loss, $\mathcal{L}_{\text{calib}}(f(x),y)$, can be defined as a squared loss: $\mathcal{L}_{\text{calib}}(\hat{\boldsymbol{p}},\boldsymbol{y})=\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}.$
@@
$\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\hat{p}_{ik})^{2}\right]-\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{ik}-\kappa(\hat{p}_{ik}))^{2}\right]$ (89)
@@
$\begin{aligned} &\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}\left((y_{ik}-\hat{p}_{ik})^{2}-\left((y_{ik}-\hat{p}_{ik})^{2}-2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})+\eta(\hat{p}_{ik})^{2}\right)\right)\right]\\ &=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})-\eta(\hat{p}_{ik})^{2}\right]\end{aligned}$
@@
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}}^{\gamma,\lambda})$ $=\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}})+\lambda\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{calib}})$ (91)
@@
When $\hat{p}_{ik}>y_{ik}$, $\eta(\hat{p}_{ik})>0$. In this case, the term $2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})$ is negative, and $\eta(\hat{p}_{ik})^{2}<0$, which means that the calibration term overall is negative. This reduction in the post-processing gap brings the prediction closer to the true label, reducing the pGap.
@@
When $\hat{p}_{ik}<y_{ik}$, $\eta(\hat{p}_{ik})>0$, and thus $2(y_{ik}-\hat{p}_{ik})\eta(\hat{p}_{ik})$ is positive. This positive term increases the predicted value, reducing the gap between $\hat{p}_{ik}$ and $y_{ik}$. Additionally, the term $\eta(\hat{p}_{ik})^{2}$ is still negative but smaller in magnitude, ensuring that the gap is reduced.
@@
$\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{FCL}})\leq\text{pGap}_{\mathcal{D}}(\mathcal{L}_{\text{focal}}).$ (92)
A.15 Ablation Study: L1 vs. L2 Loss
We did not present results for L1 or other $L_{p}$ losses in the main text because our preliminary experiments indicated that they perform worse than L2 in terms of calibration. Table 8 shows example comparisons (test error and ECE) using Mean Absolute Error (L1) vs. Mean Squared Error (L2) on CIFAR-10 and CIFAR-100. We observe that L2 consistently achieves lower ECE, so we adopt L2 (MSE) for our focal calibration loss. As shown, L2 (MSE) outperforms L1 (MAE) in terms of ECE across different architectures, which motivated our choice of L2 in all experiments.
Table 8: Comparing L1 (MAE) vs. L2 (MSE) for calibration on CIFAR-10 and CIFAR-100.
Dataset Model Loss Test Error (%) ECE (%)
CIFAR-100 ResNet-110 L1 22.38 4.45
ResNet-110 L2 21.02 4.19
WideResNet L1 19.99 4.82
WideResNet L2 19.81 2.22
DenseNet-121 L1 21.66 3.42
DenseNet-121 L2 21.80 1.31
ResNet-50 L1 21.64 5.65
ResNet-50 L2 21.81 3.71
CIFAR-10 ResNet-110 L1 5.09 1.46
ResNet-110 L2 4.75 0.95
WideResNet L1 4.12 2.44
WideResNet L2 4.16 0.92
DenseNet-121 L1 5.37 4.00
DenseNet-121 L2 4.59 0.66
ResNet-50 L1 5.08 1.78
ResNet-50 L2 4.56 0.76
A.16 CheXNet: Interpretability and Per-Class AUROC Results
This section provides an extended discussion on how FCL improves interpretability and performance in medical imaging tasks, specifically for Chest X-ray anomaly detection (CheXNet).
Interpretability with Grad-CAM Heatmaps. Grad-CAM heatmaps offer a widely accepted visualization tool for highlighting the most influential regions of a model’s activation. By adopting FCL in training, medical practitioners gain a more intuitive understanding of the model’s decision-making process, thereby enhancing trust and transparency. As illustrated in Fig. 1 (in the main text), FCL produces clearer and more precise anomaly-localization maps compared to baseline methods. This improvement not only has practical utility in diagnostic tasks but also provides stronger interpretability, crucial for the stringent reliability standards in medical applications. Consistent with Section 5.3.1, FCL achieves superior calibration metrics while delivering finer visual explanations—bridging the gap between purely numerical evaluations and real-world clinical needs.
Training Procedure. We followed the training procedure outlined by Rajpurkar et al. (Rajpurkar, 2017), including the ten-crop augmentation for both validation and test phases. Table 9 compares per-class AUROC scores for CheXNet using BCE vs. our FCL (along with earlier baselines). Note that FCL consistently yields strong performance gains, particularly in terms of average AUROC and reliability scores (ECE, MCE).
Table 9: AUROC comparison (%) over previous SOTA results on the ChestX-ray14 dataset. ECE* and MCE† are shown in parentheses.
Pathology Wang et al.  (Wang et al., 2017) Yao et al.  (Yao et al., 2017) CheXNet (BCE)  (Rajpurkar, 2017) FCL (Ours)
Atelectasis 0.716 0.772 0.8094 0.8332
Cardiomegaly 0.807 0.904 0.9248 0.9156
Effusion 0.784 0.859 0.8638 0.8862
Infiltration 0.609 0.695 0.7345 0.7151
Mass 0.706 0.792 0.8676 0.8701
Nodule 0.671 0.717 0.7802 0.8017
Pneumonia 0.633 0.713 0.7680 0.7813
Pneumothorax 0.806 0.841 0.8887 0.8866
Consolidation 0.708 0.788 0.7901 0.8192
Edema 0.835 0.882 0.8878 0.8996
Emphysema 0.815 0.829 0.9371 0.9417
Fibrosis 0.769 0.767 0.8047 0.8559
Pleural Thickening 0.708 0.765 0.8062 0.7958
Hernia 0.767 0.914 0.9164 0.9370
Average 0.742 (/) 0.798 (/) 0.8440 (0.1738*, 0.6134†) 0.8527 (0.1292*, 0.3248†)
Overall, FCL consistently enhances both calibration quality and interpretability in CheXNet. These improvements align with the high standards of reliability and transparency demanded in clinical practice, making FCL particularly suitable for real-world medical diagnostic workflows.
Figure 8: Reliability Diagram of Various Methods before Temperature Scaling (CIFAR10, ResNet-50, 100 - 300 Epochs). The top/middle/bottom row corresponds to the results at 100/200/300 epochs. Each row shows the performance of Cross Entropy, FLSD-53, Dual Focal Loss, and Focal Calibration Loss.
A.17 Reliability Diagram and ROC Diagram
Figure 9: Reliability Diagram of Various Methods before Temperature Scaling(CIFAR10, ResNet-110, 300 Epochs).